OpenStudy (anonymous):

How to muiltiply (x-1/2(1-sqrt5i))(x-1/2(1+sqrt5i))

OpenStudy (dumbcow):

(1-sqrt5i)(1+sqrt5i) = 1 -5i^2 = 1+5 = 6 (remember i^2 = -1) --> (x-1)^2 / 24

OpenStudy (anonymous):

What did you do with 1/2 , maybe i should say the problem use the given zero to find all the zeros of the function they give me 1/2(1-sqrt 5i) meaning i also have 1/2(1+sqrt5i) cause there pairs when i put them as factors i put X- in front of them and muiltiply to get somehting to divide the original function by

OpenStudy (dumbcow):

oh i read it wrong its (x-(1/2)(1-sqrt5i) ??

OpenStudy (anonymous):

Yes and then +sqrt5i

OpenStudy (dumbcow):

ok you are right the other zero would be 1/2(1 +sqrt5i) do you need the quadratic function? you could multiply them together or just find it this way: x = 1/2(1-sqrt5i) Now work backwards until you get rid of "i" and right side is zero --> 2x = 1-sqrt5i 2x -1 = -sqrt5i (2x-1)^2 = -5 (2x-1)^2 +5 = 0

OpenStudy (dumbcow):

wait do they give you a function f(x) originally

OpenStudy (anonymous):

Mhm and is that only finding one like i know i need to muiltyply both those zeros to get my number too divide by the function is 8x^3-14x^2+18x-9

OpenStudy (dumbcow):

ok i see (x - 1/2(1-sqrt5i))(x -1/2(1+sqrt5i)) =x^2 -1/2x +xsqrt5i/2 -1/2x -xsqrt5i/2 +1/4(6) = x^2 -x +3/2

OpenStudy (dumbcow):

then after dividing you should get: 8x-6 = 0 solve to get 3rd zero

OpenStudy (anonymous):

Thankk you!!! Soo much